OpenStudy (anonymous):

Which coordinate pair is in the solution set for y < 1 − 6x? a.) (-1, 1) b.) (1, −1) c.) (1, 0) d.) (0, 1) Can someone help me with this.

OpenStudy (jango_in_dtown):

y<1-6x implies y+6x<1

OpenStudy (jango_in_dtown):

a) x=-1, y=1 now put the values of x and y and you have y+6x=1+6(-1)=1-6=-5<1

OpenStudy (jango_in_dtown):

so (-1,1) is a solution

OpenStudy (jango_in_dtown):

b)x=1 y=-1 so y+6x=-1+6=5 which is not less then 1 so (1,-1) is not solution

OpenStudy (jango_in_dtown):

c)x=1 y=0 so y+6x=0+6=6 is not less than 1 so again (1,0) is not a solution

OpenStudy (jango_in_dtown):

d) x=0 y=1 so y+6x=1 is not less than 1 so (0,1) is also not possible hence option a) is only correct

OpenStudy (jango_in_dtown):

@espeletachristina

OpenStudy (anonymous):

Thank you so much you are the greatest, I understood what you were talking about thanks a lot.

OpenStudy (jango_in_dtown):

n.p.:)